40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 27 20:28:10 2013, manasa, Update, Green Locking, ALS servo OLTF.pdfphase_badget_xarm_ALS.pdf
    Reply  Tue Oct 1 17:14:53 2013, masayuki, Update, Green Locking, ALS servo filters modified OLTF.pdffilter_change.pdf
Message ID: 9171     Entry time: Fri Sep 27 20:28:10 2013     Reply to this: 9183
Author: manasa 
Type: Update 
Category: Green Locking 
Subject: ALS servo 

[Masayuki, Manasa]

I. ALS servo loops
After fixing things with the phase tracking loop, we checked if things were good with the ALS servo loops.
We measured the OLTF of the X and Y arm ALS servo loops. In both cases the phase margin was ~20 degrees. There was no room to set enough phase margin. So we looked at the servo filters. We tried to modify the filters so that we could bring enough phase margin, but could not get at it. So we put back the old filters as they were.

 attachment1: OLTF of the ALS XARM and YARM control loops

attachment2: Current phase budget. FM4 and FM10 are the boost filters.

II. ALS in-loop noise
Also, I found that the overall noise of the ALS servo has gone up by about two orders of magnitude (in Hz/rtHz) over the whole range of frequencies for both the arms from the last time the measurements were made. I suspect this could be from some change in the calibration factor. Did anybody touch things around that could have caused this? Or can somebody recollect any changes that I made in the past which might have affected the calibration? Anyways, I will do the calibration again.

 

 

Attachment 1: OLTF.pdf  10 kB  Uploaded Fri Sep 27 21:36:37 2013  | Hide | Hide all
OLTF.pdf
Attachment 2: phase_badget_xarm_ALS.pdf  13 kB  Uploaded Fri Sep 27 21:37:11 2013  | Hide | Hide all
phase_badget_xarm_ALS.pdf
ELOG V3.1.3-